OpenStudy (anonymous):

When three identical resistors are connected in series, their combined resistance is 6 ohms. What is their combined resistance when they are connected in parallel?

OpenStudy (nikvist):

\[R_S=3R\quad\Rightarrow\quad R=R_S/3\]\[R_P=R/3=R_S/9\approx 0.67\Omega\]

OpenStudy (anonymous):

 i think the answer is 3/2 ohm